# Seele: A Unified Acceleration Framework for Real-Time Gaussian Splatting

### v2

@@ 7.3 Ablation Study
MiniSplatting | 27.23 | 0.814 | 0.222 | 71.31 | 797.96 | 145.7 | 23.18 | 0.837 | 0.187 | 143.27 | 329.69 | 83.3 | 30.04 | 0.908 | 0.243 | 120.01 | 382.86 | 154.9 |
MiniSplatting+Opti. | 27.24 | 0.814 | 0.223 | 83.16 | 669.35 | 145.7 | 23.19 | 0.837 | 0.188 | 162.29 | 277.99 | 83.3 | 30.03 | 0.908 | 0.243 | 146.24 | 302.23 | 154.9 |
MiniSplatting+Opti.+HP | 27.70 | 0.823 | 0.213 | 102.95 | 421.39 | 106.5 | 23.75 | 0.846 | 0.181 | 210.31 | 169.92 | 60.1 | 30.02 | 0.908 | 0.242 | 155.67 | 197.19 | 129.6 |
MiniSplatting+Opti.+CR | 27.22 | 0.814 | 0.222 | 100.14 | 608.65 | 145.7 | 23.18 | 0.837 | 0.186 | 192.47 | 257.53 | 83.3 | 30.04 | 0.908 | 0.243 | 169.15 | 295.35 | 154.9 |
Seele + MiniSplatting | 27.70 | 0.822 | 0.212 | 131.62 | 436.41 | 106.5 | 23.74 | 0.846 | 0.179 | 268.03 | 172.40 | 60.1 | 30.02 | 0.908 | 0.242 | 200.62 | 224.58 | 129.6 |
LightGaussian | 27.44 | 0.807 | 0.235 | 30.89 | 1533.50 | 59.4 | 23.82 | 0.842 | 0.189 | 65.60 | 699.71 | 33.8 | 29.74 | 0.901 | 0.250 | 43.17 | 1097.82 | 52.4 |
LightGaussian+Opti. | 27.42 | 0.807 | 0.236 | 34.48 | 1365.02 | 59.4 | 23.83 | 0.841 | 0.190 | 72.46 | 620.48 | 33.8 | 29.72 | 0.900 | 0.251 | 48.41 | 980.20 | 52.4 |
LightGaussian+Opti.+HP | 27.64 | 0.811 | 0.230 | 70.55 | 600.20 | 39.1 | 24.04 | 0.847 | 0.182 | 172.00 | 261.86 | 20.7 | 29.74 | 0.901 | 0.250 | 126.65 | 306.88 | 41.4 |
LightGaussian+Opti.+CR | 27.43 | 0.807 | 0.235 | 40.15 | 1127.65 | 59.4 | 23.81 | 0.842 | 0.188 | 85.36 | 525.97 | 33.8 | 29.74 | 0.901 | 0.250 | 58.59 | 783.34 | 52.4 |
Seele + LightGaussian | 27.56 | 0.810 | 0.229 | 76.36 | 589.85 | 39.1 | 23.91 | 0.846 | 0.181 | 183.86 | 259.78 | 20.7 | 29.73 | 0.900 | 0.249 | 131.16 | 324.25 | 41.4 |
@@
Seele: this variant is the full-fledged algorithm including all optimizations proposed in this paper.
Tbl. 3 presents the results of our ablation study on 3DGS [27], MiniSplatting [16] and LightGaussian [15]. Across all datasets, our code optimization (+opti.) achieves a 1.1 speedup. Building on this, +Opti.+HP and +Opti.+CR yield additional 2.8 and 1.3 speedups, respectively. When all optimizations are combined, Seele achieves an overall 3.3 speedup. All model weight savings come from HP, as the other two optimizations do not contribute to the model reduction. Note that, Seele may increase the overall instruction count but improves the overall speedup due to the reductions on warp divergence.
In terms of the rendering quality, the additional optimizations proposed in Sec. 7.1 do not change the accuracy. The HP technique from Sec. 4, combined with fine-tuning as described in Sec. 6, further enhances rendering quality across all three quality metrics. On average, HP improves the rendering quality by 0.23 on PSNR. The effect of CR from Sec. 5 on quality is minimal, 0.03 in PSNR.
@@
Next, we present the results with and without the integrated fine-tuning, as described in Sec. 6. Here, we focus on the results of the Mip-NeRF360 dataset with 3DGS. Similar trends are observed across other datasets in the supplementary. As shown in Tbl. 4, incorporating fine-tuning results in significant improvements across all quality metrics. At the same time, fine-tuning does not affect performance, with both FPS and GPU memory consumption remaining relatively unchanged.
@@ 7.4 Sensitivity Study
Fig. 8 shows how rendering quality and performance vary with two key hyperparameters, the number of clusters and the number of neighboring clusters during rendering (), in Sec. 4. Performance numbers are normalized to our default setting of 24 clusters with 4 neighbors (including itself). As the number of neighboring clusters increases, speedup decreases, while accuracy initially improves. However, too many clusters, e.g., 8 clusters with 8 neighbors, degrades accuracy. This degradation is because each Gaussian is responsible for too many view directions, negatively impacting rendering quality.
Fig. 12 shows example results from our sensitivity study on rendering quality and performance to the number of clusters and cluster neighbors. All results are generated by partitioning the entire scene with 24 clusters while varying the number of neighboring clusters. We find that most artifacts come from uncovered regions, primarily due to an insufficient number of neighboring clusters. Please see details highlighted in red bounding boxes. The lack of neighboring clusters results in black ghosting shadows (Fig. 12).
@@
Fig. 9 illustrates the sensitivity of both the rendering quality and performance to the pixel group size using Mip-NeRF360 [13], as defined in Sec. 5. Other datasets show a similar trend. As the pixel group size increases, rendering quality initially decreases slightly, but the drop becomes more significant once the group size exceeds . We also present examples of rendered images with different group sizes in Fig. 9(b). As the group size increases to , artifacts begin to appear in the high-frequency regions. On the performance side, speedup increases sublinearly at first, but it almost plateaus once the group size reaches . Thus, we ultimately select .
Fig. 13 presents the qualitative results analyzing the sensitivity of both rendering quality and performance to pixel group size. One potential limitation of this technique is that as the window size increases, artifacts such as aliasing may appear in the rendered images. For instance, when the window size reaches , noticeable aliasing effects can be observed, as shown on the right side of Fig. 13.
8 Limination and Discussion
Limitation.
This paper presents Seele, a unified framework that achieves speedup across GPUs. Nevertheless, we acknowledge several limitations of this work:
Scene-Specific Configuration in Hybrid Preprocessing: In this work, we adopt a uniform configuration for clustering in hybrid preprocessing. We believe that scene-specific configurations could further enhance rendering performance and quality. However, manually tuning the number of clusters and neighboring clusters is cumbersome with a human in the loop. An interesting future work would be designing an automatic algorithm to optimize these parameters for each scene.
Aliasing in Contribution-Aware Rasterization: We observe that increasing the window size to introduces aliasing artifacts in high-frequency regions when applying contribution-aware rasterization. While our current approach balances speedup and quality, further investigation is needed to develop techniques that mitigate aliasing while maintaining efficiency.
Meanwhile, we acknowledge several works that leverage GPU fixed-function units to accelerate alpha blending using GPU RT Core [38], GPU raster [12], or WebGL [11]. Our approach is orthogonal to these methods, and we are excited to explore how our techniques can be integrated with these optimizations to further enhance rendering performance.
Looking Forward.
Recent neural rendering has changed the landscape of real-time rendering. Gaussian splatting emerges as a new rendering primitive that offers unprecedented realism using the conventional rasterization pipeline. This paper proposes two principled and GPU-oriented optimizations that drastically improve the general Gaussian splatting pipelines by understanding and leveraging the GPUs’ characteristics. We believe this work could be the first step to understanding the fundamental limitations of GPUs in supporting Gaussian splatting. The result would be valuable for rethinking brand-new architectures tailored to this future technology.

### v3

@@ 4.3 Ablation Study
Seele: this variant is the full-fledged algorithm.
Tbl. 4 presents the results of our ablation study on 3DGS. Across all datasets, our code optimization (+Opti.) achieves a 1.1 speedup. Building on this, +Opti.+HP and +Opti.+CR yield additional 2.8 and 1.3 speedups, respectively. When all optimizations are combined, Seele achieves an overall 3.2 speedup. All model weight savings come from HP, as the other two optimizations do not contribute to the model reduction. Note that, Seele may increase the overall instruction count but improves the overall speedup due to the reductions on warp divergence.
In terms of the rendering quality, the additional optimizations proposed in Sec. 4.1 do not change the accuracy. The HP technique from Sec. 3.3, combined with fine-tuning as described in Sec. 3.5, further enhances rendering quality across all three quality metrics. On average, HP improves the rendering quality by 0.23 on PSNR. The effect of CR from Sec. 3.4 on quality is minimal, 0.03 in PSNR.
@@
| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | |||
| w/ Loss | w/o Loss | w/ Loss | w/o Loss | w/ Loss | w/o Loss | |
| 3DGS | 0.058 | 0.060 | 0.067 | 0.074 | 0.042 | 0.043 |
@@
Next, we present the results with and without the integrated fine-tuning, as described in Sec. 3.5. Here, we focus on the results of the Mip-NeRF360 dataset with 3DGS. Similar trends are observed across other datasets in the supplementary. Tbl. 5 shows that fine-tuning results in significant improvements. Meanwhile, fine-tuning does not affect performance for both FPS and GPU memory consumption. Tbl. 6 also shows the effectiveness of , finetuning with our proposed loss term can improve the rendering quality.
@@ 4.4 Sensitivity Study
Fig. 6 shows how rendering quality and performance vary with two key hyperparameters, the number of clusters and the number of neighboring clusters during rendering (), in Sec. 3.3. Performance numbers are normalized to our default setting of 24 clusters with 4 neighbors (including itself). As the number of neighboring clusters increases, speedup decreases, while accuracy initially improves. However, too many clusters, e.g., 8 clusters with 8 neighbors, degrades accuracy. This degradation is because each Gaussian is responsible for too many view directions, negatively impacting rendering quality.
@@
Fig. 8 illustrates the sensitivity of both the rendering quality and performance to the pixel group size using Mip-NeRF360 [barron2021mip], as defined in Sec. 3.4. Other datasets show a similar trend. As the pixel group size increases, rendering quality initially decreases slightly, but the drop becomes more significant once the group size exceeds . By trading off between quality and performance, we ultimately select the pixel group size of .
Number of Contributors.
Fig. 8 shows the sensitivity of the rendering quality and performance to the number of primary contributors in the clustering strategy (Sec. 3.3). The rendering quality is not significantly affected by the number of contributors, while the performance gradually decreases as the number of contributors increases.
5 Conclusion
Recent neural rendering has transformed the landscape of real-time rendering. 3DGS emerges as a new rendering primitive that offers unprecedented realism using the conventional rasterization pipeline. This paper proposes two principled and GPU-oriented optimizations that drastically improve the general Gaussian splatting pipelines by understanding and leveraging the GPUs’ characteristics. We believe that this work could be the first step to understanding the fundamental limitations of GPUs in supporting 3DGS. The result would be valuable to rethink the next architectures tailored to this future technique.